# Low-Rank Continual Personalization of Diffusion Models

### v1

@@ 5 Conclusions
In this work, we analyze the continual adaptation capabilities of a DreamBooth method coupled with the Low-Rank Adaptation (LoRA) technique. We evaluate three different approaches to merging and initialization of the individual adapters. We show that naïve continual training of a LoRA leads to catastrophic forgetting, while other techniques can mitigate this issue, which originates from conflicts in adapters that are sequentially trained.
6 Acknowledgements
This research was supported by NCN OPUS grant no. 2023/51/B/ST6/03004. The computing resources were provided by PL-Grid Infrastructure grant no. PLG/2024/017176.
@@ A.1 Datasets
DreamBooth To fine-tune the model on personalized objects, the DreamBooth [24] dataset is used, which contains images of 30 objects (9 of them being dogs and cats, and 21 being other objects) belonging to 15 different classes. Each object is represented by several images (from 4 up to 6), which represent it in either a different environment or from a different camera angle. Due to these quantity limits and dataset set-up, we use the same dataset of reference images for both evaluation and fine-tuning.
UnlearnCanvas For style-related tasks, we use UnlearnCanvas collection [36], which contains 20 objects (each with 20 images) in 60 different styles. When evaluating methods for the style tasks, we select $10$ images representing $10$ different objects in the same style but the ones that were not used during training.
We generated datasets for our experiments by randomly selecting classes and objects. For object-related tasks, we selected 10 classes, each represented by a unique object and paired with a unique token. For style-based tasks, we selected 10 distinct styles, each paired with one object and five corresponding images. Object names, classes, and styles were uniquely selected to avoid repetition. Selected classes are presented in Table 3.
@@ A.2 Finetuning
We train the model on $T=10$ tasks. For each object task, we use distinct DreamBooth tokens, while for styles, we retain the real names. The learning rate is set to 1e-5, with 850 train steps and a batch size of 1. During fine-tuning, we don’t leverage the prior preservation regularization [24].
In the original paper introducing LoRA [12], the authors investigate how both the rank and the target matrices of the transformer’s self-attention module, to which the LoRA is applied, influence method efficiency. They show that, while the impact of rank $r$ is small, it is recommended to apply the LoRA to $W^{Q}$, $W^{K}$, $W^{V}$, and $W^{O}$. In our experiments, we similarly apply LoRA to these targets within the self-attention weights of the U-Net and set $r=64$.
All four methods for low-rank continual personalization methods were evaluated with two seeds (00, $5$) and four task ordering seeds (00, $5$, $10$, $42$).
@@ A.3 Evaluation metrics
To evaluate fine-tuned model generations, we use image alignment metrics to calculate the cosine similarity between embeddings of generated and reference images. For subject alignment, we use features extracted from the DINO [3] and CLIP [23] models. When assessing style, we continue using DINO but replace CLIP with CSD [31], as CLIP has demonstrated a bias towards image content, making it unsuitable for style-related tasks. When sampling from the text-to-image diffusion model, we generate $N=8$ images from each of $P=5$ prompts and calculate the mean cosine similarity of models’ embeddings between each of $N\cdot P=40$ samples and the object/style reference images. When generating an image, we do it with $50$ denoising diffusion steps.
@@
Appendix B Additional results
For style personalization, we also observe in Fig 3 (right) the origin of each method forgetting, which is the magnitude of conflicts introduced to adapters by fine-tuning on the next tasks. We see that both methods that achieve the lowest forgetting in Fig. 3 (left) - magnitude-based merging and merge & initialization - exhibit the least increasing ratio of vector norms over time. Note that for the $9$-th task, we observe plot degradation that comes from calculating conflicts according to adapters that were trained from the base model without including adapters from tasks trained later.
Refer to caption Refer to caption
Figure 3: Performance on the first task (left) and the ratio of conflicting vector norm to introduced adapter parameter (right) over continual fine-tuning on next tasks for the style personalization.
The model generally achieves its highest performance on each task immediately after fine-tuning, but this performance tends to degrade as new tasks are added. Despite this decline, our evaluated techniques can retain knowledge from previous tasks. As presented in Fig. 67 for object personalization and Fig. 89 for style personalization, we can observe that the merge & orthogonal initialization method exhibits the highest plasticity in all metrics for both objects and styles. For some cases (tasks 1 and 2 in merge & initialization method in Fig. 6), we can observe backward transfer, a situation where the performance on the previous task improves thanks to the adaptation to the new task.
For measuring $t$-th task adapter’s degradation with the next tasks incoming, we leverage two vectors of parameters: (1) task vector representing parameters introduced to the model only by fine-tuning it on that one task - calculated by subtraction of $t-1$-th adapter from the $t$-th adapter - and (2) next $m$ tasks’ vector denoting parameters introduced to the model starting from $t+1$-th task till the $t+m$-th task - calculated by subtraction of $t+m$-th adapter from the $t$-th adapter - and zero its parameters that keep the same sign as the corresponding ones in task vector. Finally, we determine the task degradation by the ratio of the norm of (2) the conflicted next tasks’ vector to the norm of (1) the task vector. We average the ratios over all the adapters, not only the first one.
@@ Appendix C Prompts for styles
Due to the complex nature of style personalization, this task is characterized by the easy overwriting of previous styles when new ones come adapted by the Dreambooth method. This fact is extremely harmful especially in the context of continued learning, where it implies catastrophic forgetting of the model. In our experiments, we note that a significant problem in the context of personalization of styles is the way they are expressed in the form of a prompt. To this end, we conduct how different forms of prompts affect the average score, average forgetting, and task scores of a diffusion model learning next styles sequentially.

### v3

@@ 5 Conclusions
In this work, we analyze the continual adaptation capabilities of a DreamBooth method coupled with the Low-Rank Adaptation (LoRA) technique. We evaluate three different approaches to merging and initialization of the individual adapters. We show that naïve continual training of a LoRA leads to catastrophic forgetting, while other techniques can mitigate this issue.
Acknowledgments
The project was funded by the National Science Centre, Poland, grants no: 2023/51/B/ST6/03004 and 2022/45/B/ST6/02817. The computational resources were provided by PLGrid grant no. PLG/2024/017114 and no. PLG/2024/017266.
Appendix
@@ A.1 Datasets
DreamBooth. To fine-tune the model on personalized objects, the DreamBooth (Ruiz et al., 2023) dataset is used, which contains images of 30 objects (9 of them being dogs and cats, and 21 being other objects) belonging to 15 different classes. Each object is represented by several images (from 4 up to 6), which represent it in either a different environment or from a different camera angle. Due to the quantity limits, we follow Ruiz et al. (2023) and use the same dataset of images for both evaluation and fine-tuning.
UnlearnCanvas. For style-related tasks, we use the UnlearnCanvas collection (Zhang et al., 2024), which contains 20 objects (each with 20 images) in 60 different styles. When evaluating methods for the style tasks, we select $10$ images representing $10$ different objects in the same style but the ones that were not used during training.
We generate datasets for our experiments by randomly selecting classes and objects. For object-related tasks, we select 10 classes, each represented by a unique object and paired with a unique token. For style-based tasks, we select 10 distinct styles, each paired with one object and five corresponding images. Object names, classes, and styles were uniquely selected to avoid repetition. Selected classes are presented in Table 3.
@@ A.2 Finetuning
We train the model on $T=10$ tasks. For each object task, we use distinct DreamBooth tokens, while for styles, we retain the real names. The learning rate is set to 1e-5, with 850 train steps and a batch size of 1. During fine-tuning, we don’t leverage the prior preservation regularization Ruiz et al. (2023).
In the original work introducing LoRA (Hu et al., 2021), the authors investigate how both the rank and the target matrices of the transformer’s self-attention module, to which the LoRA is applied, influence method efficiency. They show that, while the impact of rank $r$ is small, it is recommended to apply the LoRA to $W^{Q}$, $W^{K}$, $W^{V}$, and $W^{O}$. In our experiments, we similarly apply LoRA to these targets within both the self-attention and cross-attention layers of the U-Net and set $r=64$.
To ensure the robustness of our experiments, all four methods for low-rank continual personalization methods were evaluated with two seeds (00, $5$) and four task ordering seeds (00, $5$, $10$, $42$).
@@ A.3 Evaluation metrics
To evaluate fine-tuned model generations, we use image alignment metrics to calculate the cosine similarity between embeddings of generated and reference images. For subject alignment, we use features extracted from the DINO Caron et al. (2021) and CLIP Radford et al. (2021) models. When assessing style, we continue using DINO but replace CLIP with CSD Somepalli et al. (2024a), as CLIP has demonstrated a bias towards image content, making it unsuitable for style-related tasks. When sampling from the text-to-image diffusion model, we generate $N=8$ images from each of $P=5$ prompts and calculate the mean cosine similarity of models’ embeddings between each of $N\cdot P=40$ samples and the object/style reference images. When generating an image, we do it with $50$ denoising diffusion steps.
@@
Appendix B Style personalization results
The model generally achieves its highest performance on each task immediately after fine-tuning, but this performance tends to degrade as new tasks are added. Despite this decline, our evaluated techniques can retain knowledge from previous tasks. As presented in Figures 5 and 6 for object personalization and Figures 7 and 8 for style personalization, we can observe that the merge & orthogonal initialization method exhibits the highest plasticity in all metrics for both objects and styles. For some cases (tasks 1 and 2 in merge & initialization method in Figure 5), we can observe backward transfer, a situation where the performance on the previous task improves thanks to the adaptation to the new task.
@@ Appendix C Prompts for styles
Due to the complex nature of style personalization, this task is characterized by the easy overwriting of previous styles when new ones come adapted by the Dreambooth method. This fact is extremely harmful especially in the context of continued learning, where it implies catastrophic forgetting of the model. In our experiments, we note that a significant problem in the context of personalization of styles is the way they are expressed in the form of a prompt. To this end, we investigate how different forms of prompts affect the average score, average forgetting, and task scores of a diffusion model that learns the next styles sequentially.